# Is the Faber-Krahn inequality true for the Stokes operator?

### v1

@@ 3.1 Preliminaries
where $j_{1,1}$ is the first positive root of $J_{1}$, the Bessel function of the first kind of order 1. Finally, $E_{1}({\mathbb{B}}_{2})$ is spanned by the (eigenfunction
${\mathbf{u}}_{{\mathbb{B}}_{2}}:=\frac{J_{1}(j_{1,1}r)}{\sqrt{\pi}j_{1,1}|J_{0}(j_{1,1})|}\begin{pmatrix}-\sin\theta\\ \cos\theta\end{pmatrix}.$
and the pressure is constant.
@@ A Hilbert basis of eigenfunctions of the Dirichlet-Stokes operator is given, in polar coordinates, by
$\phi_{0,k}(r,\theta)=\frac{-J_{0}^{\prime}(\sqrt{\lambda_{0,k}}r)}{\sqrt{\pi}\sqrt{\lambda_{0,k}}|J_{0}(\sqrt{\lambda_{0,k}})|}\begin{pmatrix}-\sin\theta\\ \cos\theta\end{pmatrix},$ (11)
@@
$\begin{split}\phi_{j,k,m}(r,\theta)\ =\ &\frac{J_{j}(\sqrt{\lambda_{j,k}}r)-J_{j}(\sqrt{\lambda_{j,k}})r^{j}}{\lambda_{j,k}|J_{j}(\sqrt{\lambda_{j,k}})|r}j(-1)^{m+1}Y_{j,m}(\theta)\begin{pmatrix}\cos\theta\\ \sin\theta\end{pmatrix}\\ &+\frac{-\sqrt{\lambda_{j,k}}J_{j}^{\prime}(\sqrt{\lambda_{j,k}}r)+jJ_{j}(\sqrt{\lambda_{j,k}})r^{j-1}}{\lambda_{j,k}|J_{j}(\sqrt{\lambda_{j,k}})|}Y_{j,m+1}(\theta)\begin{pmatrix}-\sin\theta\\ \cos\theta\end{pmatrix}\end{split}$ (12)
for $j\in\mathbb{Z}^{*}$, $k\in\mathbb{N}^{*}$ and $m=1,2$
[…]
@@
[Regarding Lemmata 1-2] Lemmata 1-2 are essentially contained in [31, 36], which contain explicit computations of the eigen-elements of the Dirichlet-Stokes operator in special geometries. While we refer to these articles for detailed computations, let us briefly indicate how they might be derived in a straightforward manner: consider $({\mathbf{u}}_{k},\lambda_{k}({\mathbb{B}}_{2}))$ an eigenpair of the Dirichlet-Stokes operator. As $\nabla\cdot{\mathbf{u}}_{k}=0$, there exists a scalar function $\psi_{k}$ such that
@@
For any $\mathscr{C}^{2}$ domain ${\Omega}$ such that $\lambda_{1}({\Omega})$ is simple, the eigenvalue mapping ${\Omega}\mapsto\lambda_{1}({\Omega})$ is twice shape differentiable at $\Omega$ in the following sense: for any $\Phi\in\mathscr{X}$, the map $f_{\Phi}:t\mapsto\lambda_{1}\left((\mathrm{Id}+t\Phi){\Omega}\right)$ is twice differentiable at $t=0$. We will use the notations
@@
Similarly, the mapping ${\Omega}\mapsto{\mathbf{u}}_{\Omega}$ is twice shape-differentiable at $\Omega$, where ${\mathbf{u}}_{\Omega}$ is the first Dirichlet-Stokes normalized eigenfunction of ${\Omega}$, in the sense that the mapping $g_{\Phi}:t\mapsto{\mathbf{u}}_{(\mathrm{Id}+t\Phi)\Omega}$ is twice differentiable at $t=0$. We let ${\mathbf{u}}_{\Phi}^{\prime}$ be its derivative at $t=0$.
For a detailed introduction to the Hadamard shape calculus, we refer to [30, Chapter 5] and to [26]. To proceed with the proof of Theorem 2, we need tractable expressions for the first and second-order shape derivatives of the eigenvalue at a domain ${\Omega}$.
@@
That ${\mathbf{u}}_{\Phi}^{\prime}$ solves (14) is a standard consequence of general formulae for the shape differentiation of Dirichlet boundary value problem, and we refer to [30, Chapter 5] for the detailed computations. Once (14) is established, it suffices to multiply the equation by ${\mathbf{u}}_{\Omega}$ and to integrate by parts to obtain
$\langle d\lambda_{1}({\Omega}),\Phi\rangle\int_{\Omega}\|{\mathbf{u}}_{\Omega}\|^{2}=-\int_{\partial{\Omega}}\|({\nabla}{\mathbf{u}}_{\Omega})\nu\|^{2}\langle\Phi,\nu\rangle.$
As
$\int_{\Omega}\|{\mathbf{u}}_{\Omega}\|^{2}=1,$
the conclusion follows.
To obtain the expression for $\langle d^{2}\lambda_{1}({\Omega})\Phi,\Phi\rangle$ for normal vector fields, it suffices to apply the Hadamard formula for integrals on variable boundaries [30, Proposition 5.4.18]. This yields
$\langle d^{2}\lambda_{1}({\Omega})\Phi,\Phi\rangle$ $=-2\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,{\nabla}{\mathbf{u}}_{\Omega}\nu\rangle\langle\Phi,\nu\rangle$
@@
$\langle d^{2}\lambda_{1}({\Omega})\Phi,\Phi\rangle$ $=-2\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,{\nabla}{\mathbf{u}}_{\Omega}\nu\rangle\langle\Phi,\nu\rangle$
$-\int_{\partial{\Omega}}\left(H\|({\nabla}{\mathbf{u}}_{\Omega})\nu\|^{2}+\frac{\partial||({\nabla}{\mathbf{u}}_{\Omega})\nu||^{2}}{\partial\nu}\right)\langle\Phi,\nu\rangle^{2}$
$=-2\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,{\nabla}{\mathbf{u}}_{\Omega}\nu\rangle\langle\Phi,\nu\rangle$
$+\int_{\partial{\Omega}}H\|({\nabla}{\mathbf{u}}_{\Omega})\nu\|^{2}\langle\Phi,\nu\rangle^{2}-2\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Omega})\nu,{\nabla}p_{\Omega}\rangle\langle\Phi,\nu\rangle^{2}$
@@ 3.2 The ball is a critical point for Problem (5) (in dimension two)
Fix $\Phi\in\mathscr{X}$. The derivative of the volume $\mathrm{Vol}:{\Omega}\mapsto|{\Omega}|$ at ${\mathbb{B}}_{2}$ in the direction $\Phi$ is given by
@@
However, this relation holds because of the explicit expression given in Lemma 1. ∎
@@ 3.3 The ball satisfies strong second-order conditions for Problem (5)
Using Proposition 1 (and noting first that $p_{{\mathbb{B}}_{2}}$, the pressure, is constant here and, second that $H=1$) we obtain
$\langle d^{2}\mathcal{F}({\mathbb{B}}_{2})\Phi,\Phi\rangle$ $=\lambda_{1}({\mathbb{B}}_{2})\int_{\partial{\mathbb{B}}_{2}}\varphi^{2}-2\int_{\partial{\Omega}}\varphi\int_{\partial{\mathbb{B}}_{2}}\|({\nabla}{\mathbf{u}}_{{\mathbb{B}}_{2}})\nu\|^{2}\varphi$
$+2|{\mathbb{B}}_{2}|\cdot\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,{\mathbf{u}}_{\Phi}^{\prime}\rangle+|{\mathbb{B}}_{2}|\cdot\int_{\partial{\mathbb{B}}_{2}}\|({\nabla}{\mathbf{u}}_{{\mathbb{B}}_{2}})\nu\|^{2}\varphi^{2}.$
@@
$\langle d^{2}\mathcal{F}({\mathbb{B}}_{2})\Phi,\Phi\rangle$ $=2\lambda_{1}({\mathbb{B}}_{2})\left(\int_{\partial{\mathbb{B}}_{2}}\varphi^{2}\right)-\frac{2\lambda_{1}({\mathbb{B}}_{2})}{|{\mathbb{B}}_{2}|}\left(\int_{\partial{\mathbb{B}}_{2}}\varphi\right)^{2}+2|{\mathbb{B}}_{2}|\cdot\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,{\mathbf{u}}_{\Phi}^{\prime}\rangle$
@@
$=-\int_{\partial{{\mathbb{B}}_{2}}}p^{\prime}\langle(\nabla{{\mathbf{u}}_{{\mathbb{B}}_{2}}}\nu),{\nu}\rangle\varphi$
@@
Write ${{\mathbf{u}}}_{{\mathbb{B}}_{2}}=(u_{1},u_{2})^{\top}$, and ${{\mathbf{u}}_{\Phi}}^{\prime}=(u_{\Phi,1}^{\prime},u_{\Phi,2}^{\prime})^{\top}$. With a slight abuse of notation, we write, in polar coordinates, ${{\mathbf{u}}}_{{\mathbb{B}}_{2}}(x)={{\mathbf{u}}}(r,\theta)=\phi_{0,1}(r,\theta){\vec{e}_{r}}$, where $\phi_{0,1}$ is given by (11). Similarly, as we know that we can assume the perturbation $\Phi$ to be normal, we identify $\Phi$ with $\varphi$, and we decompose $\varphi:\partial{\mathbb{B}}^{2}\to\mathbb{R}$ as a Fourier series by considering the two sequences $(\alpha_{n})_{n\in\mathbb{N}}$ and $(\beta_{n})_{n\in\mathbb{N}}$ in $\ell^{2}(\mathbb{N})$ such that
@@
$\Delta^{2}\psi_{\Phi}^{\prime}+\lambda_{1}({\mathbb{B}}_{2})\Delta\psi_{\Phi}^{\prime}=0\quad\text{in }B_{2}.$
@@
$\Delta\psi_{\Phi}^{\prime}+\lambda_{1}({\mathbb{B}}_{2})\psi_{\Phi}^{\prime}=\sum_{n=0}^{+\infty}(c_{n}\cos(n\theta)+d_{n}\sin(n\theta))r^{n}.$
Solving this equation in $\psi_{\Phi}^{\prime}$ and using the Bessel functions, we thus deduce that there exist two sequences $(a_{n})_{n\in\mathbb{N}}\,,(b_{n})_{n\in\mathbb{N}}\in\ell^{2}(\mathbb{N})$ such that
@@
$u_{1}^{\prime}=\frac{\partial\psi_{\Phi}^{\prime}}{\partial x_{2}}=-\frac{\partial u_{1}}{\partial\nu}\varphi\quad\text{on }\partial{\mathbb{B}}_{2}\qquad\text{and}\qquad u_{2}^{\prime}=-\frac{\partial_{\Phi}^{\prime}}{\partial x_{1}}=-\frac{\partial u_{2}}{\partial\nu}\varphi\quad\text{on }\partial{\mathbb{B}}_{2}.$
Recall that
@@
$\begin{pmatrix}u_{1}^{\prime}\\ u_{2}^{\prime}\end{pmatrix}=\frac{j_{1,1}}{\sqrt{\pi}}\begin{pmatrix}\sin\theta\\ -\cos\theta\end{pmatrix}\sum_{n=0}^{+\infty}(\alpha_{n}\cos(n\theta)+\beta_{n}\sin(n\theta))\quad\text{on }\partial{\mathbb{B}}_{2}.$
Using that $u_{1}^{\prime}=\partial\psi_{\Phi}^{\prime}/\partial x_{2}$, it follows that
$u_{1}^{\prime}=\sum_{n=0}^{+\infty}\sin\theta j_{1,1}J_{n}^{\prime}(j_{1,1})(a_{n}\cos(n\theta)+b_{n}\sin(n\theta))+n\sin\theta(c_{n}\cos(n\theta)+d_{n}\sin(n\theta))$
$\displaystyle+\sum_{n=0}^{+\infty}n\cos\theta J_{n}(j_{1,1})(b_{n}\cos(n\theta)-a_{n}\sin(n\theta))+n\cos\theta(d_{n}\cos(n\theta)-c_{n}\sin(n\theta))\quad\text{on }\partial{\mathbb{B}}_{2}.$
Identifying the cosine and sine terms in the relation above yields
  • in front of the constant term:
    $j_{1,1}J_{1}^{\prime}(j_{1,1})b_{1}+d_{1}=-\frac{j_{1,1}}{\sqrt{\pi}}\beta_{1},$
  • in front of the term $\cos\theta$:

### v2

@@ 3.1 Preliminaries
where $j_{1,1}$ is the first positive root of $J_{1}$, the Bessel function of the first kind of order 1. Finally, $E_{1}({\mathbb{B}}_{2})$ is spanned by the eigenfunction
${\mathbf{u}}_{{\mathbb{B}}_{2}}:=\frac{J_{1}(j_{1,1}r)}{\sqrt{\pi}|J_{0}(j_{1,1})|}\begin{pmatrix}-\sin\theta\\ \cos\theta\end{pmatrix},$
normalised in $L^{2}({\mathbb{B}}_{2})$, and the associated pressure is constant.
Remark 8 (A useful computation).
We single out the following computations, which will prove useful when studying the local optimality of the ball. Let us set $c=1/(\sqrt{\pi}|J_{0}(j_{1,1})|)$, $r=\sqrt{x^{2}+y^{2}}$ and $f(r)=J_{1}(j_{1,1}r)/r$. For $r>0$, one easily computes
$\nabla{\mathbf{u}}_{{\mathbb{B}}_{2}}=c\begin{pmatrix}-\frac{xy}{r}f^{\prime}(r)&-f(r)-\frac{y^{2}}{r}f^{\prime}(r)\\ f(r)+\frac{x^{2}}{r}f^{\prime}(r)&\frac{xy}{r}f^{\prime}(r)\end{pmatrix}$
so that
$\nabla{\mathbf{u}}_{{\mathbb{B}}_{2}}=cf^{\prime}(1)\begin{pmatrix}-\cos\theta\sin\theta&-\sin^{2}\theta\\ \cos^{2}(\theta)&\cos\theta\sin\theta\end{pmatrix}\qquad\text{on }\partial{\Omega}.$
In particular, observe that
$({\nabla}{\mathbf{u}}_{{\mathbb{B}}_{2}})\nu=cf^{\prime}(1)\begin{pmatrix}-\sin(\theta)\\ \cos(\theta)\end{pmatrix}.$
@@ A Hilbert basis of eigenfunctions of the Dirichlet-Stokes operator is given, in polar coordinates, by
$\phi_{0,k}(r,\theta)=\frac{-J_{0}^{\prime}(\sqrt{\lambda_{0,k}}r)}{\sqrt{\pi}|J_{0}(\sqrt{\lambda_{0,k}})|}\begin{pmatrix}-\sin\theta\\ \cos\theta\end{pmatrix},$ (10)
@@
$\begin{split}\phi_{j,k,m}(r,\theta)\ =\ &\frac{J_{j}(\sqrt{\lambda_{j,k}}r)-J_{j}(\sqrt{\lambda_{j,k}})r^{j}}{\sqrt{\lambda_{j,k}}|J_{j}(\sqrt{\lambda_{j,k}})|r}j(-1)^{m+1}Y_{j,m}(\theta)\begin{pmatrix}\cos\theta\\ \sin\theta\end{pmatrix}\\ &+\frac{-\sqrt{\lambda_{j,k}}J_{j}^{\prime}(\sqrt{\lambda_{j,k}}r)+jJ_{j}(\sqrt{\lambda_{j,k}})r^{j-1}}{\sqrt{\lambda_{j,k}}|J_{j}(\sqrt{\lambda_{j,k}})|}Y_{j,m+1}(\theta)\begin{pmatrix}-\sin\theta\\ \cos\theta\end{pmatrix}\end{split}$ (11)
for $j\in\mathbb{Z}^{*}$, $k\in\mathbb{N}^{*}$ and $m=1,2$.
Here, $(r,\theta)$ are the usual polar coordinates (see [36, 41]). The functions $Y_{j,m}(\theta)$ are defined by $Y_{j,1}(\theta)=\frac{1}{\sqrt{\pi}}\cos(j\theta)$ and $Y_{j,2}(\theta)=\frac{1}{\sqrt{\pi}}\sin(j\theta)$, with the agreement that $Y_{j,3}=Y_{j,1}$, and $J_{j}$ is the Bessel function of the first kind of order $j$. Denoting by $j_{j,k}>0$ the $k^{\textrm{th}}$ positive zero of $J_{j}$, the eigenvalues of the Dirichlet-Stokes operator are the doubly indexed sequence $(\lambda_{j,k})_{j\in\mathbb{Z},k\in\mathbb{N}^{*}}$, where $\lambda_{j,k}=j_{|j|+1,k}^{2}$ is of multiplicity $1$ if $j=0$, and $2$ if $j\neq 0$.
@@
[Regarding Lemmata 1-2] Lemmata 1-2 are essentially contained in [36, 41], which provide explicit computations of the eigen-elements of the Dirichlet-Stokes operator in special geometries. While we refer to these articles for detailed computations, let us briefly indicate how they might be derived in a straightforward manner: consider $({\mathbf{u}}_{k},\lambda_{k}({\mathbb{B}}_{2}))$ an eigenpair of the Dirichlet-Stokes operator. As $\nabla\cdot{\mathbf{u}}_{k}=0$, there exists a scalar function $\psi_{k}$ such that
@@
For any $\mathscr{C}^{2}$ domain ${\Omega}$ such that $\lambda_{1}({\Omega})$ is simple, the eigenvalue mapping ${\Omega}\mapsto\lambda_{1}({\Omega})$ is twice differentiable at $\Omega$ in the following sense: for any $\Phi\in\mathscr{X}$, the map $f_{\Phi}:t\mapsto\lambda_{1}\left((\mathrm{Id}+t\Phi){\Omega}\right)$ is twice differentiable at $t=0$. We will use the notations
@@
Similarly, the mapping ${\Omega}\mapsto{\mathbf{u}}_{\Omega}$ is twice differentiable at $\Omega$, where ${\mathbf{u}}_{\Omega}$ is the first Dirichlet-Stokes normalized eigenfunction of ${\Omega}$, in the sense that the mapping $g_{\Phi}:t\mapsto{\mathbf{u}}_{(\mathrm{Id}+t\Phi)\Omega}$ is twice differentiable at $t=0$. We let ${\mathbf{u}}_{\Phi}^{\prime}$ be its derivative at $t=0$.
For a detailed introduction to the Hadamard shape calculus, we refer to [34, Chapter 5] and to [30]. To proceed with the proof of Theorem 2, we need tractable expressions for the first and second-order shape derivatives of the eigenvalue at a domain ${\Omega}$.
@@
That ${\mathbf{u}}_{\Phi}^{\prime}$ solves (13) is a standard consequence of general formulae for the shape differentiation of Dirichlet boundary value problem, and we refer to [34, Chapter 5] for the detailed computations. To derive the expression of $\langle d\lambda_{1}({\Omega}),\Phi\rangle$, multiply (13) by $u_{\Omega}$ to obtain:
$\langle d\lambda_{1}({\Omega}),\Phi\rangle$ $=\langle d\lambda_{1}({\Omega}),\Phi\rangle\int_{\Omega}\|{\mathbf{u}}_{\Omega}\|^{2}\quad\left(\text{ since }\int_{\Omega}\|{\mathbf{u}}_{\Omega}\|^{2}=1\right)$
$=\int_{\Omega}{\nabla}{\mathbf{u}}_{\Phi}^{\prime}\colon{\nabla}{\mathbf{u}}_{\Omega}-\underbrace{\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,{\nabla}{\mathbf{u}}_{\Omega}\rangle}_{=0\text{ since }{\mathbf{u}}_{\Omega}=0\text{ on }\partial{\Omega}}+\underbrace{\int_{\Omega}\langle{\nabla}p^{\prime},{\mathbf{u}}_{\Omega}\rangle}_{=0\text{ since }{\nabla}\cdot{\mathbf{u}}_{\Omega}=0}-\lambda_{1}({\Omega})\underbrace{\int_{\Omega}\langle{\mathbf{u}}_{\Phi}^{\prime},{\mathbf{u}}_{\Omega}\rangle}_{=0\text{ since }\int_{\Omega}\langle{\mathbf{u}}_{\Omega}\,,{\mathbf{u}}_{\Phi}^{\prime}\rangle=0}$
$=-\int_{\Omega}\langle\Delta{\mathbf{u}}_{\Omega}\,,{\mathbf{u}}_{\Phi}^{\prime}\rangle+\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Omega})\nu,{\mathbf{u}}_{\Phi}^{\prime}\rangle$
$=-\underbrace{\int_{\Omega}\langle{\nabla}p_{\Omega},{\mathbf{u}}_{\Phi}^{\prime}\rangle}_{=0\text{ since }{\nabla}\cdot{\mathbf{u}}_{\Phi}^{\prime}=0}+\lambda_{1}({\Omega})\underbrace{\int_{\Omega}\langle{\mathbf{u}}_{\Phi}^{\prime},{\mathbf{u}}_{\Omega}\rangle}_{=0\text{ since }\int_{\Omega}\langle{\mathbf{u}}_{\Omega}\,,{\mathbf{u}}_{\Phi}^{\prime}\rangle=0}+\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Omega})\nu,{\mathbf{u}}_{\Phi}^{\prime}\rangle$
$=-\int_{\partial{\Omega}}\|({\nabla}{\mathbf{u}}_{\Omega})\nu\|^{2}\langle\Phi,\nu\rangle.$
To obtain the expression for $\langle d^{2}\lambda_{1}({\Omega})\Phi,\Phi\rangle$ for normal vector fields, it suffices to apply the Hadamard formula for integrals on variable boundaries [34, Proposition 5.4.18]. This yields
$\langle d^{2}\lambda_{1}({\Omega})\Phi,\Phi\rangle$ $=-2\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,({\nabla}{\mathbf{u}}_{\Omega})\nu\rangle\langle\Phi,\nu\rangle$
@@
In particular, this implies, taking into account that $\Delta_{\tau}{\mathbf{u}}_{\Omega}=0\,,{\mathbf{u}}_{\Omega}=0$ on $\partial{\Omega}$,
$\frac{\partial(({\nabla}{\mathbf{u}}_{\Omega})\nu)}{\partial\nu}=\Delta{\mathbf{u}}_{\Omega}-\Delta_{\tau}{\mathbf{u}}_{\Omega}-H({\nabla}{\mathbf{u}}_{\Omega})\nu=-\lambda_{1}({\Omega}){\mathbf{u}}_{\Omega}-{\nabla}p_{\Omega}-H({\nabla}{\mathbf{u}}_{\Omega})\nu=-{\nabla}p_{\Omega}-H({\nabla}{\mathbf{u}}_{\Omega})\nu.$
@@
$\frac{\partial\|({\nabla}{\mathbf{u}}_{\Omega})\nu\|^{2}}{\partial\nu}=2\left\langle({\nabla}{\mathbf{u}}_{\Omega})\nu,\frac{\partial(({\nabla}{\mathbf{u}}_{\Omega})\nu)}{\partial\nu}\right\rangle=-2\langle{\nabla}p_{\Omega}\,,({\nabla}{\mathbf{u}}_{\Omega})\nu\rangle-2H\|({\nabla}{\mathbf{u}}_{\Omega})\nu\|^{2},$ (16)
so that
$H\|({\nabla}{\mathbf{u}}_{\Omega})\nu\|^{2}+\frac{\partial||({\nabla}{\mathbf{u}}_{\Omega})\nu||^{2}}{\partial\nu}=-H\|({\nabla}{\mathbf{u}}_{\Omega})\nu\|^{2}-2\langle{\nabla}p_{\Omega}\,,({\nabla}{\mathbf{u}}_{\Omega})\nu\rangle$
and, finally, since $({\nabla}{\mathbf{u}}_{\Omega})\nu\langle\Phi,\nu\rangle=-{\mathbf{u}}_{\Phi}^{\prime}$ on $\partial{\Omega}$,
$\langle d^{2}\lambda_{1}({\Omega})\Phi,\Phi\rangle$ $=-2\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,({\nabla}{\mathbf{u}}_{\Omega})\nu\rangle\langle\Phi,\nu\rangle-\int_{\partial{\Omega}}\left(H\|({\nabla}{\mathbf{u}}_{\Omega})\nu\|^{2}+\frac{\partial||({\nabla}{\mathbf{u}}_{\Omega})\nu||^{2}}{\partial\nu}\right)\langle\Phi,\nu\rangle^{2}$
@@ 3.2 The ball is a critical point for Problem (5) (in dimension two)
Fix $\Phi\in\mathscr{X}$. Recall that $\mathscr{X}$ has been introduced in Lemma 3. The derivative of the volume $\mathrm{Vol}:{\Omega}\mapsto|{\Omega}|$ at ${\mathbb{B}}_{2}$ in the direction $\Phi$ is given by
@@
Let us now prove that (18) indeed holds. Recall from Remark 8 that if we introduce $f(r):=J_{1}(j_{1,1}r)$ then we have, on $\partial{\Omega}$,
$({\nabla}{\mathbf{u}}_{{\mathbb{B}}_{2}})\nu=cf^{\prime}(1)\begin{pmatrix}-\sin(\theta)\\ \cos(\theta)\end{pmatrix},\text{ with }c=\frac{1}{\sqrt{\pi}|J_{0}(j_{1,1})|}\text{ whence }\|({\nabla}{\mathbf{u}}_{{\mathbb{B}}_{2}})\nu\|^{2}=\frac{f^{\prime}(1)^{2}}{\pi J_{0}(j_{1,1})^{2}}.$
Since $J_{1}^{\prime}(x)=J_{0}(x)-J_{1}(x)/x$ for any $x>0$, we further derive
$f^{\prime}(1)=j_{1,1}J_{1}^{\prime}(j_{1,1})=j_{1,1}J_{0}(j_{1,1}).$
Consequently,
$\|({\nabla}{\mathbf{u}}_{{\mathbb{B}}_{2}})\nu\|^{2}=\frac{f^{\prime}(1)^{2}}{\pi J_{0}(j_{1,1})^{2}}=\frac{j_{1,1}^{2}}{\pi}=\frac{j_{1,1}^{2}}{|{\mathbb{B}}_{2}|}.$
The conclusion follows. ∎
@@ 3.3 The ball satisfies strong second-order conditions for Problem (5)
Using Proposition 1, the fact that $p_{{\mathbb{B}}_{2}}$, the pressure, is constant (see Lemma 1) and that $H=1$ we obtain
$\langle d^{2}\mathcal{F}({\mathbb{B}}_{2})\Phi,\Phi\rangle$ $=\lambda_{1}({\mathbb{B}}_{2})\int_{\partial{\mathbb{B}}_{2}}\varphi^{2}-2\int_{\partial{\mathbb{B}}_{2}}\varphi\int_{\partial{\mathbb{B}}_{2}}\|({\nabla}{\mathbf{u}}_{{\mathbb{B}}_{2}})\nu\|^{2}\varphi$
$+2|{\mathbb{B}}_{2}|\cdot\int_{\partial{\mathbb{B}}_{2}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,{\mathbf{u}}_{\Phi}^{\prime}\rangle+|{\mathbb{B}}_{2}|\cdot\int_{\partial{\mathbb{B}}_{2}}\|({\nabla}{\mathbf{u}}_{{\mathbb{B}}_{2}})\nu\|^{2}\varphi^{2}.$
@@
$\langle d^{2}\mathcal{F}({\mathbb{B}}_{2})\Phi,\Phi\rangle$ $=2\lambda_{1}({\mathbb{B}}_{2})\left(\int_{\partial{\mathbb{B}}_{2}}\varphi^{2}\right)-\frac{2\lambda_{1}({\mathbb{B}}_{2})}{|{\mathbb{B}}_{2}|}\left(\int_{\partial{\mathbb{B}}_{2}}\varphi\right)^{2}+2|{\mathbb{B}}_{2}|\cdot\int_{\partial{\mathbb{B}}_{2}}\langle({\nabla}{\mathbf{u}}_{\Phi}^{\prime})\nu,{\mathbf{u}}_{\Phi}^{\prime}\rangle$
@@
$=-\int_{\partial{{\mathbb{B}}_{2}}}p^{\prime}\langle(\nabla{{\mathbf{u}}_{{\mathbb{B}}_{2}}})\nu,{\nu}\rangle\varphi$
@@
Write ${{\mathbf{u}}}_{{\mathbb{B}}_{2}}=(u_{1},u_{2})^{\top}$, and ${\mathbf{u}}_{\Phi}^{\prime}\color[rgb]{0,0,0}=(u_{\Phi,1}^{\prime},u_{\Phi,2}^{\prime})^{\top}$. With a slight abuse of notation, we write, in polar coordinates, ${{\mathbf{u}}}_{{\mathbb{B}}_{2}}(x)={{\mathbf{u}}}(r,\theta)=\phi_{0,1}(r,\theta){\vec{e}_{\theta}}$, where $\phi_{0,1}$ is given by (10) and ${\vec{e}_{\theta}}=\begin{pmatrix}-\sin\theta\\ \cos\theta\end{pmatrix}$ Similarly, as we know that we can assume the perturbation $\Phi$ to be normal, we identify $\Phi$ with $\varphi$, and we decompose $\varphi:\partial{\mathbb{B}}_{2}\to\mathbb{R}$ as a Fourier series by considering the two sequences $(\alpha_{n})_{n\in\mathbb{N}}$ and $(\beta_{n})_{n\in\mathbb{N}}$ in $\ell^{2}(\mathbb{N})$ such that
@@
$\Delta^{2}\psi_{\Phi}^{\prime}+\lambda_{1}({\mathbb{B}}_{2})\Delta\psi_{\Phi}^{\prime}=0\quad\text{in }{\mathbb{B}}_{2}\color[rgb]{0,0,0}.$
@@
$\Delta\psi_{\Phi}^{\prime}+\lambda_{1}({\mathbb{B}}_{2})\psi_{\Phi}^{\prime}={\lambda_{1}({\mathbb{B}}_{2})}\sum_{n=0}^{+\infty}(c_{n}\cos(n\theta)+d_{n}\sin(n\theta))r^{n}.$
Observe that the function $G$ given in polar coordinates by $G:(r,\theta)\mapsto\sum_{n=0}^{+\infty}(c_{n}\cos(n\theta)+d_{n}\sin(n\theta))r^{n}$ is a particular solution of this equation in ${\mathbb{B}}_{2}$. As the function $z$ given in polar coordinates by $z:(r,\theta)\mapsto\psi_{\Phi}^{\prime}(r,\theta)-G(r,\theta)$ solves $\Delta z+\lambda_{1}({\mathbb{B}}_{2})z=0$, a separation of variables yields the existence of two sequences $(a_{n})_{n\in\mathbb{N}}\,,(b_{n})_{n\in\mathbb{N}}\in\ell^{2}(\mathbb{N})$ such that
@@
$u_{1}^{\prime}=-\frac{\partial\psi_{\Phi}^{\prime}}{\partial y}=-\frac{\partial u_{1}}{\partial\nu}\varphi\quad\text{on }\partial{\mathbb{B}}_{2}\qquad\text{and}\qquad u_{2}^{\prime}=\frac{\partial\psi_{\Phi}^{\prime}}{\partial x}=-\frac{\partial u_{2}}{\partial\nu}\varphi\quad\text{on }\partial{\mathbb{B}}_{2}.$
As $J_{1}^{\prime}(x)=J_{0}(x)-J_{1}(x)/x$ for any $x>0$, we deduce that $J_{1}^{\prime}(j_{1,1})=J_{0}(j_{1,1})<0$. From Remark 8, one gets
@@
$\begin{pmatrix}u_{1}^{\prime}\\ u_{2}^{\prime}\end{pmatrix}=\frac{J_{0}(j_{1,1})}{|J_{0}(j_{1,1})|}\cdot\frac{j_{1,1}}{\sqrt{\pi}}\begin{pmatrix}-\sin\theta\\ \cos\theta\end{pmatrix}\varphi(\theta)=-\frac{j_{1,1}}{\sqrt{\pi}}\begin{pmatrix}-\sin\theta\\ \cos\theta\end{pmatrix}\varphi(\theta)\quad\text{on }\partial{\mathbb{B}}_{2}.$
Using that $u_{1}^{\prime}=-\partial\psi_{\Phi}^{\prime}/\partial y$ and $u_{2}^{\prime}=\partial\psi_{\Phi}^{\prime}/\partial x$, it follows that